The market price of a share of preferred stock is $23.71 and the dividend is $2.40. What discount rate did the market use to value the stock?

Suppose the risk-free rate is 1.43% and an analyst assumes a market risk premium of 7.51%. Firm A just paid a dividend of $1.49 per share. The analyst estimates the β of Firm A to be 1.22 and estimates the dividend growth rate to be 4.61% forever. Firm A has 286.00 million shares outstanding. Firm B just paid a dividend of $1.55 per share. The analyst estimates the β of Firm B to be 0.85 and believes that dividends will grow at 2.70% forever. Firm B has 200.00 million shares outstanding. What is the value of Firm A?

A firm will pay a dividend of $2.67 next year. The dividend is expected to grow at a constant rate of 2.32% forever and the required rate of return is 10.49%. What is the value of the stock?

A firm just paid a dividend of $3.75. The dividend is expected to grow at a constant rate of 2.08% forever and the required rate of return is 13.72%. What is the value of the stock?
Answer #1

Discount rate = (Preferred dividend / price) * 100

Discount rate = (2.4 / 23.71) * 100

Discount rate = 10.12%
Answer #2

1/ 2.4/ 23.7 = 0.10 =10%

answered by: Ayyas
Answer #3

3/ Div0 (1+g)/r-g = 15(1.0232)/ 0.1049-0.0232= 33.43

answered by: Ayyas
Answer #4

4/ Div0 (1+g)/r-g = 3.75(1.0208)/ 0.1372-0.0208 = 32.88

answered by: Ayyas